### v2

# Scaling Properties of Diffusion Models for Perceptual Tasks

Abstract
In this paper, we argue that iterative computation with diffusion models offers a powerful paradigm for not only generation but also visual perception tasks. We unify tasks such as depth estimation, optical flow, and amodal segmentation under the framework of image-to-image translation, and show how diffusion models benefit from scaling training and test-time compute for these perceptual tasks. Through a careful analysis of these scaling properties, we formulate compute-optimal training and inference recipes to scale diffusion models for visual perception tasks. Our models achieve competitive performance to state-of-the-art methods using significantly less data and compute. We release code and models at scaling-diffusion-perception.github.io.
**footnotetext: Equal Contribution
1 Introduction
Diffusion models have emerged as powerful techniques for generating images and videos, while showing excellent scaling behaviors. However, diffusion models can also be utilized for downstream prediction for inverse vision problems. In this paper, we present a unified framework to perform depth estimation, optical flow estimation, and amodal segmentation with a single diffusion model, as illustrated in Figure 1.
Previous works such as Marigold (Ke et al., 2024), FlowDiffuser (Luo et al., 2024), and pix2gestalt (Ozguroglu et al., 2024) demonstrate the potential of repurposing image diffusion models for various inverse vision tasks individually. Building on this foundation, we perform an extensive empirical study, establishing scaling power laws for depth estimation, and display their transferability to other perceptual tasks. Using insights from these scaling laws, we formulate compute-optimal recipes for diffusion training and inference. We are the first to show that efficiently scaling compute for diffusion models leads to significant performance gains in downstream perceptual tasks.
Recent works have also focused on scaling test-time compute to enhance the capabilities of modern LLMs, as demonstrated by OpenAI’s o1 model (OpenAI, 2024). As Noam Brown noted, “It turned out that having a bot think for just 20 seconds in a hand of poker got the same boosting performance as scaling up the model by 100,000x and training it for 100,000 times longer.” In our experiments, we realize a similar trade off between allocating more compute during training versus test-time for diffusion models.
We scale test-time compute by exploiting the iterative and stochastic nature of diffusion to increase the number of denoising steps. By allocating more compute to early denoising steps, and ensembling multiple denoised predictions, we consistently achieve higher accuracy on these perceptual tasks. Our results provide evidence of the benefits of scaling test-time compute for inverse vision problems under constrained compute budgets, bringing a new perspective to the conventional paradigm of training-centric scaling for generative models.
Figure 1: A Unified Framework: We fine-tune a pre-trained Diffusion Model (DM), for visual perception tasks. We take a RGB image, and a conditional image (i.e. next video frame, occlusion mask, etc.), along with the noised image of the ground truth prediction. Our model generates predictions for visual tasks such as depth estimation, optical flow prediction, and amodal segmentation, based on the conditional task embedding. We train a generalist model that can perform all three tasks with exceptional performance.
2 Related Work
Generative Modeling: Generative modeling has been studied under various methods, including VAEs (Kingma, 2013), GANs (Goodfellow et al., 2014), Normalizing Flows (Rezende & Mohamed, 2015), Autoregressive models (van den Oord et al., 2016), and Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) have shown impressive scaling behaviors for many image and video generation models. Notable examples include Latent Diffusion Models (Rombach et al., 2022), which enhanced efficiency by operating in a compressed latent space, Imagen (Saharia et al., 2022), which generates samples in pixel space with increasing resolution, and Consistency Models (Song et al., 2023), which aim to accelerate sampling while maintaining generation quality. Recent methods like Rectified Flow (Liu et al., 2022) and Flow Matching (Lipman et al., 2023) employ training objectives inspired by optimal transport to model continuous vector fields that map data to target distributions, eliminating the discrete formulation of diffusion models. Rectified Flow mitigates numerical issues in training by applying flow regularization, and Flow Matching offers efficient sampling with fewer discretization artifacts, making them promising alternatives to diffusion for high-quality generation. Apart from diffusion models, Parti (Yu et al., 2022) and MARS (He et al., 2024) showcased the potential of autoregressive models for image generation, and the Muse architecture (Chang et al., 2023) introduced a masked image generation approach using transformers.
Scaling Diffusion Models: Diffusion modeling has shown impressive scaling behaviors in terms of data, model size, and compute. Latent Diffusion Models (Rombach et al., 2022) first showed that training with large-scale web datasets can achieve high quality image generation results with a U-Net backbone. DiT (Peebles & Xie, 2023) explored scaling diffusion models with the transformer architecture, presenting desirable scaling properties for class-conditional image generation. Later, Li et al.(Li et al., 2024) studied alignment scaling laws of text-to-image diffusion models. Recently, Fei et al.(Fei et al., 2024a) trained mixture-of-experts DiT models up to 16B parameters, achieving high-quality image generation results. Upcycling is another way to scale transformer models. Komatsuzaki et al. (Komatsuzaki et al., 2022) used upcycling to convert a dense transformer-based language model to a mixture-of-experts model without pre-training from scratch. Similarly, EC-DiTSun et al. (2024) explores how to exploit heterogeneous compute allocation in mixture-of-experts training for DiT models through expert-choice routing and learning to adaptively optimize the compute allocated to specific text-image data samples.
Diffusion Models for Perception Tasks: Diffusion models have also been used for various downstream visual tasks such as depth estimation (Ji et al., 2023; Duan et al., 2023; Saxena et al., 2023; 2024; Zhao et al., 2023). More recently, Marigold (Ke et al., 2024) and GeoWizard (Fu et al., 2024) displayed impressive results by repurposing pre-trained diffusion models for monocular depth estimation. Diffusion models with few modifications are used for semantic segmentation for categorical distributions (Hoogeboom et al., 2021; Brempong et al., 2022; Tan et al., 2022; Amit et al., 2021; Baranchuk et al., 2021; Wolleb et al., 2022), instance segmentation (Gu et al., 2024), and panoptic segmentation (Chen et al., 2023). Diffusion models are also used for optical flow (Luo et al., 2024; Saxena et al., 2024) and 3D understanding (Liu et al., 2023; Jain et al., 2022; Poole et al., 2022; Wang et al., 2023; Watson et al., 2022).
3 Generative Pre-Training
We first explore how to efficiently scale diffusion model pre-training. We pre-train diffusion models for class-conditional image generation using a diffusion transformer (DiT) backbone and follow the original model training recipe (Peebles & Xie, 2023).
Starting with a target RGB image $I\in\mathbb{R}^{u\times u\times 3}$, where the resolution of the image is $u\times u$, our pretrained, frozen Stable Diffusion variational autoencoder (Rombach et al., 2022) compresses the target to a latent $z_{0}\in\mathbb{R}^{w\times w\times 4}$, where $w=u/8$. Gaussian noise is added at sampled time steps to obtain a noisy target latent. Noisy samples are generated as:
$z_{t}=\sqrt{\alpha_{t}}\cdot z_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon_{t}$ (1)
for timestep $t$. The noise is distributed as $\epsilon\sim\mathcal{N}(0,I)$, $t\sim\text{Uniform}(T)$, with $T=1000$ and $\alpha_{t}:=\prod_{s=1}^{t}(1-\beta_{s})$, with $\{\beta_{1},\dots,\beta_{T}\}$ as the variance schedule of a process.
In the denoising process, the class-conditional DiT $f_{\theta}(\cdot)$, parameterized by learned parameters $\theta$, gradually removes noise from $z_{t}$ to obtain $z_{t-1}$. The parameters $\theta$ are updated by noising $z_{0}$ with sampled noise $\epsilon$ at a random timestep $t$, computing the noise estimate, and optimizing the mean squared loss between the generated noise and estimated noise in an $n$ batch size sample. We formally represent this as the following minimization problem:
$\theta^{*}=\arg\min_{\theta}\mathcal{L}_{\theta}(z_{t},\epsilon_{i})=\arg\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}(\epsilon_{i}-\hat{\epsilon}_{i})^{2},$ (2)
where $\theta^{*}$ are the DiT learned parameters and $\hat{\epsilon}_{i}$ is the DiT noise prediction for sample $i$.
3.1 Model Size
Figure 2: Scaling at Model Size: For generative pre-training of DiT models, we observe clear power law scaling behavior as we increase the model size.
We pre-train six different dense DiT models as in Table 2, increasing model size by varying the number of layers and hidden dimension size. We use Imagenet-1K (Russakovsky et al., 2015) as our pre-training dataset and train all models for 400k iterations with a fixed learning rate of $1e$-$4$ and a batch size of 256. Fig. 2 shows that larger models converge to lower loss with a clear power law scaling behavior. We show the train loss as a function of compute (in MACs), and our predictions indicate a power law relationship of $L(C)=0.23\times C^{-0.0098}$. Our pre-training experiments display the ease of scaling DiT with a small training dataset, which translates directly to efficiently scaling downstream model performance.
3.2 Mixture of Experts
We also pre-train Sparse Mixture of Experts (MoE) models (Shazeer et al., 2017), following the S/2 and L/2 model configurations in (Fei et al., 2024b). We use three different MoE configurations listed in Table 2, scaling the total parameter count by increasing hidden size, number of experts, layers, and attention heads. Each MoE block activates the top-2 experts per token and has a shared expert that is used by all tokens. To alleviate issues with expert balance, we use the proposed expert balance loss function from (Fei et al., 2024b) which distributes the load across experts more efficiently. Sparse MoE pre-training allows for a higher parameter count while increasing throughput, making it more compute efficient than training a dense DiT model of the same size. We train our DiT-MoE models with the same training recipe as the dense DiT model using ImageNet-1K. Our approach enables training DiT-MoE models to increase model capacity without increasing compute usage by another order of magnitude, which would be required to train dense models of similar sizes.
Model Params Dimension Heads Layers
a1 14.8M 256 16 12
a2 77.2M 512 16 16
a3 215M 768 16 20
a4 458M 1024 16 24
a5 1.2B 1536 16 28
a6 1.9B 1792 16 32
Table 1: Dense DiT Models: We scale dense DiT model size by increasing hidden dimension and number of layers linearly while keeping number of heads constant following (Yang et al., 2022; Touvron et al., 2023).
Model Active / Total Dim Heads Layers
S/2-8E2A 71M / 199M 384 6 12
S/2-16E2A 71M / 369M 384 6 12
L/2-8E2A 1.0B / 2.8B 1024 16 24
Table 2: MoE DiT Models: We scale the MoE DiT models by increasing dimension size, number attention heads, layers, and experts following (Fei et al., 2024b).
4 Fine-Tuning for Perceptual tasks
In this section, we explore how to scale the fine-tuning of the pre-trained DiT models to maximize performance on downstream perception tasks. During fine-tuning, we utilize the image-to-image diffusion process from  (Ke et al., 2024) and  (Brooks et al., 2023) as our training recipe. We pose all our visual tasks as conditional denoising diffusion generation. Give an RGB image $I\in\mathbb{R}^{u\times u\times 3}$ and its pair ground truth image $D\in\mathbb{R}^{u\times u\times 3}$, we first project them to the latent space, $i_{0}\in\mathbb{R}^{w\times w\times 4}$ and $d_{0}\in\mathbb{R}^{w\times w\times 4}$, respectively. We only add noise to the ground truth latent to obtain $d_{t}$ and concatenate it with the RGB latent which results in a tensor $z_{t}=\{i_{0},d_{t}\}$. The first convolutional layer of the DiT model is modified to match the doubled number of input channels, and its values are reduced by half to make sure the predictions are the same if the inputs are just RGB images (Ke et al., 2024). Finally, we perform diffusion training by denoising the ground truth image. We ablate several fine-tuning compute scaling techniques on the monocular depth estimation task and report Absolute Relative error and Delta1 error. We transfer the best configurations from the depth estimation ablation study to fine-tune for other visual perception tasks.
4.1 Effect of Model Size
We fine-tune the pre-trained a1-a6 dense models on the depth estimation task to study the effect of model size. We scale model size as shown in as described in Section 3.1. Fig. 3 shows that larger dense DiT models predictably converge to a lower fine-tuning loss, presenting a clear power law scaling behavior. We plot the train loss and validation metrics as a function of compute (in MACs). Our fine-tuned model predictions show a power law relationship in both depth Absolute Relative error and depth Delta1 error. These experiments provide strong signal on how model performance will scale as we increase fine-tuning compute by scaling model size.
Figure 3: Effect of Model Size: We fine-tune a1-a6 models on the Hypersim dataset for 30K iterations with an exponential decay learning rate schedule from $3e$-$5$ to $3e$-$7$. We observe a strong correlation between the fine-tuning loss scaling law and validation metric scaling laws.
4.2 Effect of Pre-training Compute
We also investigate the behavior of fine-tuning as we scale the number of pre-training steps for the DiT backbone. We train four models with the a4 configuration using a varied number of pre-training steps, keeping all other hyperparameters constant. We then fine-tune these four models on the same depth estimation dataset.Fig. 4 displays the power law scaling behavior of the validation metrics for depth estimation as we increase DiT pre-training steps. Our experiments show that having stronger pre-trained representations can be helpful when scaling fine-tuning compute.
Figure 4: Effect of Scaling Model Pre-training Compute on Depth Estimation: (a) Depth Absolute Relative Error vs. MACs. (b) Depth Delta1 Error vs. MACs. We pre-train four a4 models with 60K, 80K, 100K, and 120K steps. These models are then fine-tuned for 30K steps on the Hypersim depth estimation dataset. We observe a clear power law as we increase the DiT pre-training compute across depth estimation validation metrics.
4.3 Effect of Image Resolution
The sequence length of each image also affects the total compute spent during training. For each forward pass, we can scale the amount of compute used by simply increasing the resolution of the image, which will increase the number of tokens in the image embedding. By increasing the number of tokens, we can increase the amount of information the model can learn from at training time to build stronger internal representations, which can in turn improve downstream performance. We use dense DiT-XL models with resolutions of $256\times 256$ and $512\times 512$ from (Peebles & Xie, 2023) and we pre-train DiT-MoE L/2-8E2A models with $256\times 256$ and $512\times 512$ resolutions following the recipe in (Fei et al., 2024b). We then fine-tune each of these models with the corresponding resolution for the depth estimation task. Fig. 5 displays that increasing image resolution to scale fine-tuning compute can provide significant gains on downstream depth estimation performance.
Figure 5: Effect of Image Resolution. We fine-tune DiT-XL and DiT-MoE L/2 models with resolutions of $256\times 256$ and $512\times 512$. We observe a power law when increasing image resolution during training. By scaling the number of tokens per image by 4$\times$, we achieve strong performance on Depth Absolute Error, displaying the effect of increasing total dataset tokens for dense visual perception tasks such as depth estimation.
4.4 Effect of Upcycling
Sparse MoE models are efficient options for increasing the capacity of a model, but pre-training an MoE model from scratch can be expensive. One way to alleviate this issue is Sparse MoE Upcycling (Komatsuzaki et al., 2023). Upcycling converts a dense transformer checkpoint to an MoE model by copying the MLP layer in each transformer block $E$ times, where $E$ is the number of experts, and adding a learnable router module that sends each token to the top-$k$ selected experts. The outputs of the selected experts are then combined in a weighted sum at the end of each MoE block. We upcycle various dense DiT models after they are fine-tuned for depth estimation and then continue fine-tuning the upcycled model. Fig. 6 displays the scaling laws for upcycling, providing an average improvement of 5.3% on Absolute Relative Error and 8.6% on Delta1 error.
Figure 6: Effect of Upcycling. We upcycle a2, a3, and a4 models fine-tuned for depth estimation with a varying number of total/active model experts. We continue fine-tuning each upcycled model for 15K iterations on the Hypersim depth estimation dataset. We observe a clear scaling law in the validation metrics as we increase fine-tuning compute with upcycling. The upcycled models can also achieve equivalent or superior performance to our dense a5 and a6 checkpoints, each of which utilize more compute during pre-training and fine-tuning. Increasing the total model experts and total active experts can also improve the downstream performance.
5 Scaling Test-Time Compute
Scaling test-time compute has been explored for autoregressive Large Language Models (LLMs) to improve performance on long-horizon reasoning tasks (Brown et al., 2024; Snell et al., 2024; El-Refai et al., 2024; OpenAI, 2024). In this section, we show how to reliably improve diffusion model performance for perceptual tasks by scaling test-time compute. We summarize our approach in Fig. 7. We use the Stable-Diffusion VAE to encode the input image into latent space (Rombach et al., 2022). Then, we sample a target noise latent from a standard Gaussian distribution, which is iteratively denoised with DDIM (Song et al., 2021) to generate the downstream prediction.
Figure 7: Inference Scaling: Diffusion models by design allow efficient scaling of test-time compute. First, we can simply increase the number of denoising steps to increase the compute spent at inference. Since we are estimating deterministic outputs, we can then initialize multiple noise latents and ensemble the predictions to get a better estimation. Finally, we can also reallocate our test-time compute budget for low and high frequency denoising by modifying the noise variance schedule.
5.1 Effect of Scaling Inference Steps
The most natural way of scaling diffusion inference is by increasing denoising steps. Since the model is trained to denoise the input at various timesteps, we can scale the number of diffusion denoising steps at test-time to produce finer, more accurate predictions. This coarse-to-fine denoising paradigm is also reflected in the generative case, and we can take advantage of it for the discriminative case by increasing the number of denoising steps. In Fig. 8, we observe that increasing the total test-time compute by simply increasing the number of diffusion sampling steps provides substantial gains in depth estimation performance.
Figure 8: Effect of Number of Sampling Steps. (a) Delta1 Error vs. Number of Steps. (b) Absolute Relative Error vs. Number of Steps. For each model, we sample for $T\in\left[1,2,5,10,20,50,100\right]$ steps with the DDIM sampler. We show a clear power law scaling behavior in (a) and (b), displaying the effectiveness of scaling test-time compute by increasing the number of diffusion sampling steps.
5.2 Effect of Test Time Ensembling
We also explore scaling inference compute with test-time ensembling. We exploit the fact that denoising different noise latents will generate different downstream predictions. In test-time ensembling, we compute $N$ forward passes for each input sample and reduce the outputs through one of two methods. The first technique is naive ensembling where we use the pixel-wise median across all outputs as the prediction. The second technique presented in Marigold (Ke et al., 2024) is median compilation, where we collect predictions $\{\hat{{\bm{d}}_{1}},\dots,\hat{{\bm{d}}_{N}}\}$ that are affine-invariant, jointly estimate scale and shift parameters $\hat{s_{i}}$ and $\hat{t_{i}}$, and minimize the distances between each pair of scaled and shifted predictions $(\hat{{\bm{d}}^{\prime}_{i}},\hat{{\bm{d}}^{\prime}_{j}})$ where $\hat{{\bm{d}}^{\prime}}=\hat{{\bm{d}}}\times\hat{s}+\hat{t}$. For each optimization step, we take the pixel-wise median ${\bm{m}}(x,y)=\text{median}(\hat{{\bm{d}}^{\prime}_{1}(x,y)},\dots,\hat{{\bm{d}}^{\prime}_{N}(x,y)})$ to compute the merged depth ${\bm{m}}$. Since it requires no ground truth, we scale ensembling by increasing $N$ to utilize more test-time compute.
Figure 9: Effect of Test Time Ensembling. (a) Delta1 Error vs. Number of Forward Passes. (b) Absolute Relative Error vs. Number of Forward Passes. Ensembling multiple predictions from distinct noise initializations displays power law scaling behavior. We apply test-time ensembling values of $N\in\left[1,2,5,10,15,20\right]$.
5.3 Effect of Noise Variance Schedule
We can also scale test-time compute by increasing compute usage at different points of the denoising process. In diffusion noise schedulers, we can define a schedule for the variance of the Gaussian noise applied to the image over the total diffusion timesteps $T$. Tuning the noise variance schedule allows for reorganizing compute by allocating more compute to denoising steps earlier or later in the noise schedule. We experiment with three different noise level settings for DDIM: linear, scaled linear, and cosine. Cosine scheduling from (Nichol & Dhariwal, 2021) linearly declines from the middle of the corruption process, ensuring the image is not corrupted too quickly as in linear schedules. Fig. 10 shows that the cosine noise variance schedule outperforms linear schedules for DDIM on the depth estimation task under a fixed compute budget.
Figure 10: Effect of Noise Variance (Beta) Schedule. We fine-tune a4 models with three different beta schedules: linear, scaled linear, cosine. Reallocating compute with the cosine schedule to spend more time denoising at earlier timesteps significantly improved Delta1 and Absolute Relative Error rates.
6 Putting It All Together
Using the lessons from our scaling experiments on depth estimation, we train diffusion models for optical flow prediction and amodal segmentation. We show that using diffusion models while considering efficient methods to scale training and test-time compute can provide substantial performance gains on visual perception tasks, achieving improved or similar performance as current state-of-the-art techniques. Our experiments provide insight on how to efficiently apply diffusion models for these visual perception tasks under limited compute budgets. Finally, we train a unified expert model, capable of performing all three visual perception tasks previously mentioned, displaying the generalizability of our method. Our results prove the effectiveness of our training and test-time scaling strategies, removing the need to use pre-trained models trained on internet-scale datasets to enable high-quality visual perception in diffusion models. Fig. 11 displays the predicted samples from our models.
6.1 Depth Estimation
We combine our findings from the ablation studies on depth estimation to create a model with the best training and inference configurations. We train a DiT-XL model from (Peebles & Xie, 2023) on depth estimation data from Hypersim for 30K steps with a batch size of 1024, resolution of $512\times 512$, and a learning rate exponentially decaying from $1.2e$-$4$ to $1.2e$-$6$. We use median compilation ensembling with a cosine noise variance schedule. From our scaling experiments, we found the optimal configuration for inference to be 200 denoising steps with $N=5$ samples for ensembling. As shown in Table 3, our model achieves the same validation performance as Marigold on the Hypersim dataset and better performance on the ETH3D test set while being trained with lower resolution images and approximately three orders of magnitude less pre-training data and compute.
Method Hypersim ETH3D NYUv2 ScanNet DIODE
AbsRel $\downarrow$ $\delta 1\uparrow$ AbsRel $\downarrow$ $\delta 1\uparrow$ AbsRel $\downarrow$ $\delta 1\uparrow$ AbsRel $\downarrow$ $\delta 1\uparrow$ AbsRel $\downarrow$ $\delta 1\uparrow$
DiverseDepth -- -- 22.8 69.4 11.7 87.5 10.9 88.2 37.6 63.1
MiDaS -- -- 18.4 75.2 11.1 88.5 12.1 84.6 33.2 71.5
LeReS -- -- 17.1 77.7 9.0 91.6 9.1 91.7 27.1 76.6
Omnidata -- -- 16.6 77.8 7.4 94.5 7.5 93.6 33.9 74.2
HDN -- -- 12.1 83.3 6.9 94.8 8.0 93.9 24.6 78.0
DPT -- -- 7.8 94.6 9.8 90.3 8.2 93.4 18.2 75.8
Marigold 13.5 87.5 6.5 96.0 5.5 96.4 6.4 95.1 30.8 77.3
Ours 13.6 87.6 4.8 97.8 6.8 95.0 7.7 93.7 31.0 77.2
Table 3: Depth Estimation Performance Comparison on Multiple Datasets. We achieve state-of-the-art performance on the ETH3D dataset and competitive performance across all other benchmarks. Notably, we closely match the performance of Marigold across all datasets with significantly less training compute.
6.2 Optical Flow Prediction
Optical flow estimation involves predicting the motion of objects between consecutive frames in a video, represented as a dense vector field indicating pixel-wise displacement. We use a similar configuration as the depth estimation model for optical flow training. We train a DiT-XL model on the FlyingChairs dataset for 40K steps with batch size of 1024, resolution of $512\times 512$, and learning rate exponentially decaying from $1.2e$-$4$ to $1.2e$-$6$. We compare our model’s performance with other specialized optical flow prediction techniques in Table 4.
Method FlyingChairs EPE $\downarrow$
DeepFlow 3.53
FlowNetS 2.71
FlowNetS+v 2.86
FlowNetS+ft 3.04
FlowNetS+ft+v 3.03
FlowNetC 2.19
FlowNetC+v 2.61
FlowNetC+ft 2.27
FlowNetC+ft+v 2.67
Ours (w/o ensembling) 3.45
Ours (w/ ensembling) 3.08
Table 4: Optical Flow Comparison with Specialized Techniques. We evaluate our optical flow model on the FlyingChairs validation set. Our model achieves similar end-point error as specialized methods, including DeepFlow (Weinzaepfel et al., 2013) and FlowNet (Fischer et al., 2015). We train with significantly less data compared to other specialized methods, which use a several optical flow datasets. We generate predictions with and without test-time ensembling.
6.3 Amodal Segmentation
Amodal segmentation is the process of predicting the complete shape and extent of objects in an image, including the portions that are occluded or not directly visible, which can require higher-level reasoning for complex scenes. We fine-tune a DiT-XL model on the pix2gestalt dataset (Ozguroglu et al., 2024) for 6K steps with a batch size of 4096, resolution of $256\times 256$, and learning rate exponentially decaying from $1.2e$-$4$ to $1.2e$-$6$. We compare our model with other methods in Table 5.
Method COCO-A P2G MP3D
PCNet 81.35 -- --
PCNet-Sup 82.53 -- --
SAM 67.21 -- --
SD-XL Inpainting 76.52 -- --
pix2gestalt 82.9 88.7 61.5
Ours 82.9 88.6 63.9
Table 5: Amodal Segmentation Performance (mIOU) Comparison Across Different Datasets. This table compares mIOU performance across COCO-A, Pix2Gestalt, and MP3D datasets, showing the effectiveness of various methods. Our method is able to achieve competitive performance across all tasks, while training only on Pix2Gestalt.
6.4 One Model for All
We train a unified DiT-XL model for each of the different tasks. We train this model on a mixed dataset consisting of all three tasks. To train this generalist model, we modify the DiT-XL architecture by replacing the patch embedding layer with a separate PatchEmbedRouter module, which routes each VAE embedding to a specific input convolutional layer based perception task. This ensures the DiT-XL model is able to distinguish between the task-specific embeddings during fine-tuning. We use a similar training recipe as the previous experiments, using images with $512\times 512$ resolution and a learning rate exponentially decaying from $1.2e$-$4$ to $1.2e$-$6$. Then, we upcycle the fine-tuned DiT-XL checkpoint to an DiT-XL-8E2A model, and continue fine-tuning for another 4K iterations. We display the generated predictions in Fig. 11 which exemplify the generalizability and transferability of our scaling techniques across a variety of perception tasks.
Figure 11: Depth Estimation, Optical Flow Estimation, and Amodal Segmentation Examples: Each row showcases results from our models for different tasks. (a) Depth estimation, with relative scale and shift. (b) Optical flow, with scale and shift. (c) Amodal segmentation, where the model sees an RGB image and segmentation of the occluded object; the task is to predict the amodal image.
7 Conclusion
In our work, we examine the scaling properties of diffusion models for visual perception tasks. We explore various approaches to scale diffusion training, including increasing model size, mixture-of-experts models, increasing image resolution, and upcycling. We also efficiently scale test-time compute by exploiting the iterative nature of diffusion, which significantly improves downstream performance. Our experiments provide strong evidence of scaling, uncovering power laws across various training and inference scaling techniques. We hope to inspire future work in scaling training and test-time compute for iterative generative paradigms such as diffusion for perception tasks.
8 Acknowledgments
We thank Alexei Efros for helpful discussions. We also thank Xinlei Chen, Amil Dravid, Neerja Thakkar for their valuable feedback on the paper.